OpenStudy (luigi0210):

Integration by partial fractions:

OpenStudy (luigi0210):

\[\LARGE \int \frac{x^3+4}{x^2+4}~dx\]

OpenStudy (luigi0210):

\[\LARGE \int x~dx-\int\frac{4x-4}{x^2+4}~dx\] \[\LARGE \int x~dx-4\int\frac{x-1}{x^2+4}~dx\]

ganeshie8 (ganeshie8):

Looks good, you may break the second integral into two - 1) work one integral using u-substitution 2) work the other integral using trig-substitution

OpenStudy (luigi0210):

Alright, just wasn't sure if I did it right up to that point, thank you :)

random231 (random231):

yesh its right! :)

ganeshie8 (ganeshie8):

absolutely right ! you did the long division correctly to reduce the degree of numerator and everything else looks right !

ganeshie8 (ganeshie8):

\[ \int x~dx-4\int\frac{x-1}{x^2+4}~dx = \int x~dx-4\int\frac{x}{x^2+4}~dx - 4\int\frac{1}{x^2+4}~dx \]